OpenStudy (anonymous):

how do you solve these problems: 2n(squared) - 11n + 15 and -3n(squared) - 75

OpenStudy (anonymous):

are you dealing with imaginary numbers?

OpenStudy (anonymous):

no i am dealing with factoring polynomials

OpenStudy (anonymous):

so 2n^2-11n+15 = 2n^2 -6n-5n+15

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

-3n^2-75=0--------> n^2=-25

OpenStudy (anonymous):

thank you soo much!!

OpenStudy (anonymous):

you are welcome. note that for the second equation, you will have imaginary roots.

OpenStudy (radar):

Since we are tasked to factor the polynomials, take a close look at the second equation.\[-3n ^{2}-75\]

OpenStudy (radar):

It is a factor problem, extract the -3 which is present in both terms.\[-3(n ^{2}+25)\]

OpenStudy (radar):

Thats is about as far as you can go with this.

OpenStudy (anonymous):

That's true, unless the right hand side is specified in the equation.

OpenStudy (radar):

I didn't notice any but if it is zero, we would have the situation that you mentioned and probably doehrman hasn't got into complex number yet.

OpenStudy (anonymous):

yes, you are right